2. A concrete highway is built of slabs 16.0 m long (at 17 °C). How wide should the expansion cracks be (at 17 °C) between the slabs to prevent buckling if the annual extreme temperatures are −32 °C and 54 °C?(the coefficient of linear expansion of concrete is 1.20 × 10 − 5 °C-1)
Answer #1

Here,

as the concrete expands on heating

increase in temperature , T = 54 - 17 = 37 degree C

width of gaps = L * a * T

width of gaps = 16 * 1.2 *10^-5 * 37

width of gaps = 7.1 *10-3 m = 7.1 mm

the width of gap needed is 7.1 mm